OpenStudy (anonymous):

Two forces with magnitudes of 200 and 100 pounds act on an object at angles of 60° and 170° respectively. Find the direction and magnitude of these forces. Round to two decimal places in all intermediate steps and in your final answer.

OpenStudy (tkhunny):

Can you use the angles to decompose the forces to their orthogonal component forces?

OpenStudy (anonymous):

im not even sure. It is for pre calculus, So i don think that has anything to do withorthogonal component forces.

OpenStudy (tkhunny):

Oh, yes it does. 200 lbs at 60º is \(200\cdot\cos(60º)\) in the horizontal direction and \(200\cdot\sin(60º)\) in the vertical direction. Horizontal and Vertical are these orthogonal component forces. You don't have to call them that.

OpenStudy (anonymous):

okay, I understand that, but would it be the same thing as above for the 100lbs and 170 degrees?

OpenStudy (tkhunny):

Yes. You just need to be sure you know which way things go on the horizontal and vertical axes. Something pushing from 60º will push DOWN on a standard vertical axis and LEFT on a standard horizontal axis. Something pushing from 170º will push DOWN on a standard vertical axis and RIGHT on a standard horizontal axis. Draw the forces and draw a little right traingle on each. You should see it.